# MolTC: Towards Molecular Relational Modeling In Language Models

### 13 Feb 2024 v3

@@ Abstract
Molecular Relational Learning (MRL), aiming to understand interactions between molecular pairs, plays a pivotal role in advancing biochemical research. Recently, the adoption of large language models (LLMs), known for their vast knowledge repositories and advanced logical inference capabilities, has emerged as a promising way for efficient and effective MRL. Despite their potential, these methods predominantly rely on textual data, thus not fully harnessing the wealth of structural information inherent in molecular graphs. Moreover, the absence of a unified framework exacerbates the issue of insufficient data exploitation, as it hinders the sharing of interaction mechanism learned across various datasets. To address these challenges, this work proposes a novel LLM-based multi-modal framework for Molecular inTeraction modeling following Chain-of-Thought (CoT) theory, termed MolTC, which effectively integrate graphical information of two molecules in pair. For achieving a unified training paradigm, MolTC innovatively develops a Dynamic Parameter-sharing Strategy for cross-dataset information exchange. Moreover, to train this integrated framework efficiently, we introduce a Multi-hierarchical CoT theory to refine its training paradigm, and conduct a comprehensive Molecular Interactive Instructions dataset for the development of biochemical LLMs involving MRL. Our experiments, conducted across various datasets involving over 4,000,000 molecular pairs, exhibit the superiority of our method over current GNN and LLM-based baselines. Code is available at https://github.com/MangoKiller/MolTC.
@@ 1 Introduction
(2) Dataset Foundation Construction: In sight of the absence of a comprehensive MRL datasets for biochemical LLMs, we construct a Molecular inTeractive instructions dataset, termed MoT-instruction. Specifically, we first conduct eleven well-established MRL datasets across various domain, and source their detailed molecular properties from authoritative biochemical databases. Based on this, we meticulously compile these properties and empirically determine their optimal instructions. These process ensures that MoT-instructions can not only enhance the performance of our MolTC, but also contribute to the development of other biochemical LLMs involving MRL.
@@ 3 Experiment
Baseline Model CombiSolv-QM Abraham CompSol CombiSolv
@@
GNN Based CIGIN $0.077_{\pm 0.002}$ $0.176_{\pm 0.004}$ $0.314_{\pm 0.004}$ $0.607_{\pm 0.006}$ $0.197_{\pm 0.003}$ $0.349_{\pm 0.005}$ $0.288_{\pm 0.002}$ $0.664_{\pm 0.004}$
D-MPNN $0.116_{\pm 0.006}$ $0.208_{\pm 0.005}$ $0.484_{\pm 0.004}$ $0.705_{\pm 0.005}$ $0.205_{\pm 0.006}$ $0.373_{\pm 0.007}$ $0.482_{\pm 0.003}$ $0.895_{\pm 0.005}$
GEM $0.079_{\pm 0.003}$ $0.162_{\pm 0.002}$ $\underline{0.264}_{\pm 0.004}$ $0.551_{\pm 0.005}$ $0.203_{\pm 0.006}$ $0.337_{\pm 0.007}$ $0.290_{\pm 0.006}$ $0.783_{\pm 0.009}$
CGIB $\underline{0.074}_{\pm 0.004}$ $\underline{0.150}_{\pm 0.005}$ $0.267_{\pm 0.007}$ $\underline{0.547}_{\pm 0.004}$ $\underline{0.178}_{\pm 0.004}$ $\underline{0.301}_{\pm 0.003}$ $\underline{0.251}_{\pm 0.004}$ $\underline{0.658}_{\pm 0.005}$
ML Based GOVER $0.094_{\pm 0.003}$ $0.277_{\pm 0.005}$ $0.347_{\pm 0.001}$ $0.625_{\pm 0.003}$ $0.184_{\pm 0.001}$ $0.371_{\pm 0.003}$ $0.412_{\pm 0.002}$ $0.728_{\pm 0.004}$
SolvBert $0.102_{\pm 0.005}$ $0.318_{\pm 0.006}$ ${0.496}_{\pm 0.006}$ ${0.693}_{\pm 0.006}$ $0.192_{\pm 0.008}$ $0.353_{\pm 0.008}$ ${0.418}_{\pm 0.007}$ ${0.711}_{\pm 0.004}$
Uni-Mol ${0.089}_{\pm 0.006}$ ${0.214}_{\pm 0.005}$ $0.355_{\pm 0.007}$ $0.602_{\pm 0.008}$ $0.198_{\pm 0.002}$ $0.344_{\pm 0.003}$ $0.267_{\pm 0.004}$ $0.669_{\pm 0.005}$
SMD $0.107_{\pm 0.004}$ $0.341_{\pm 0.003}$ $0.400_{\pm 0.004}$ $0.646_{\pm 0.005}$ $0.199_{\pm 0.006}$ $0.348_{\pm 0.007}$ $0.657_{\pm 0.004}$ $1.023_{\pm 0.005}$
LLM Based Galactica $0.303_{\pm 0.004}$ $0.601_{\pm 0.008}$ $0.645_{\pm 0.004}$ $1.064_{\pm 0.005}$ $0.594_{\pm 0.006}$ $0.854_{\pm 0.008}$ $0.831_{\pm 0.007}$ $1.486_{\pm 0.008}$
Chem T5 $0.321_{\pm 0.006}$ $0.555_{\pm 0.008}$ $0.629_{\pm 0.006}$ $0.910_{\pm 0.005}$ $0.445_{\pm 0.008}$ $0.734_{\pm 0.009}$ $0.882_{\pm 0.007}$ $1.297_{\pm 0.007}$
MolCA $0.298_{\pm 0.004}$ $0.545_{\pm 0.007}$ $0.581_{\pm 0.007}$ $0.897_{\pm 0.008}$ $0.467_{\pm 0.006}$ $0.716_{\pm 0.005}$ $0.648_{\pm 0.006}$ $1.125_{\pm 0.007}$
MolT5 $0.214_{\pm 0.004}$ $0.339_{\pm 0.009}$ $0.549_{\pm 0.008}$ $0.832_{\pm 0.006}$ $0.476_{\pm 0.003}$ $0.695_{\pm 0.004}$ $0.652_{\pm 0.005}$ $1.124_{\pm 0.003}$
MolTC (Ours) $0.072_{\pm 0.002}$ $0.140_{\pm 0.003}$ $0.260_{\pm 0.004}$ $0.536_{\pm 0.005}$ $0.172_{\pm 0.006}$ $0.295_{\pm 0.004}$ $0.238_{\pm 0.003}$ $0.550_{\pm 0.004}$
@@ 3.1 Experimental Setting
We evaluate MolTC on several well-established downstream molecule interaction tasks involving qualitative and quantitative analysis. Here we provide a brief overview of our experimental setup. Detailed descriptions are presented in the appendix.
Datasets. We employ 11 datasets across various domains such as DDI, SSI, and CSI. Specifically, we collect Drugbank (Version 5.0.3), ZhangDDI Zhang et al. (2017), ChChMiner Zitnik et al. (2018), DeepDDI Ryu et al. (2018), TWOSIDES Tatonetti et al. (2012), Chromophore Joung et al. (2020), MNSol Marenich et al. (2020), CompSol Moine et al. (2017), Abraham Grubbs et al. (2010), CombiSolv Vermeire and Green (2021) and CombiSolv-QM Vermeire and Green (2021).
@@ 3.3 Quantitative Prediction Performance (RQ2)
Obs.3: MolTC continues to lead in quantitative analysis tasks, an area typically challenging for LLMs. Despite the strong baseline set by CGIB, characterized by low MAE and RMSE across datasets, MolTC outperforms it in every metric. For instance, it achieves a 19.32% reduction in RMSE on the CombiSolv dataset relative to CGIB. This underscores the advantage of adeptly leveraging the interaction between SMILE representations and molecular graph structures.
@@ B.1 Datasets
We employ 11 datasets across various domains such as DDI, SSI and CSI.
@@
SSI MNSol MAE $0.584\pm{0.095}$ $0.504\pm{0.038}$ $0.491\pm{0.053}$ $\underline{0.449}\pm{0.081}$ $0.377\pm{0.026}$
RMSE $1.002\pm{0.101}$ $0.973\pm{0.079}$ $0.930\pm{0.062}$ $\underline{0.858}\pm{0.069}$ $0.668\pm{0.029}$
@@ B.6 More Experimental Results
Table 4 presents the experimental results not shown in the main text due to length constraints. Note that the three datasets in the CSI domain are all derived by splitting the Chromophore dataset. As discussed in Section 3.3, for a fair comparison, we limited the input features to four-dimensional molecular information, comprising atom type, chirality tag, bond type, and bond direction. Given the difficulty of convergence for some DL-based baselines under this setting, we only showcased the performance of the LLM-based baselines. Observations from Table 4 are largely consistent with those in the main experimental section. That is, across all tasks, our MolTC outperforms the LLM-based baseline methods in a large margin.
@@ Appendix C Future Work
In this paper, we introduce a novel unified framework, leveraging LLM technology to predict molecular interactive properties. The future development directions of this project are twofold. First, there is an emphasis on expanding its application scope, for instance, applying it to downstream tasks such as few-shot learning. Second, we aim to enhance its capabilities by incorporating technologies like graph explainability Fang et al. (2023a, b), graph sampling Wang et al. (2022, 2023), and spatio-temporal modeling Xia et al. (2023); Wu et al. (2023), making it more comprehensive or enabling it to process multiple inputs simultaneously, instead of just two.

### 21 Feb 2024 v5

@@ Abstract
Molecular Relational Learning (MRL), aiming to understand interactions between molecular pairs, plays a pivotal role in advancing biochemical research. Recently, the adoption of large language models (LLMs), known for their vast knowledge repositories and advanced logical inference capabilities, has emerged as a promising way for efficient and effective MRL. Despite their potential, these methods predominantly rely on textual data, thus not fully harnessing the wealth of structural information inherent in molecular graphs. Moreover, the absence of a unified framework exacerbates the issue of insufficient data exploitation, as it hinders the sharing of interaction mechanism learned across various datasets. To address these challenges, this work proposes a novel LLM-based multi-modal framework for Molecular inTeraction modeling following Chain-of-Thought (CoT) theory, termed MolTC, which effectively integrate graphical information of two molecules in pair. For achieving a unified training paradigm, MolTC innovatively develops a Dynamic Parameter-sharing Strategy for cross-dataset information exchange. Moreover, to train this integrated framework efficiently, we introduce a Multi-hierarchical CoT theory to refine its training paradigm, and conduct a comprehensive Molecular Interactive Instructions dataset for the development of biochemical LLMs involving MRL. Our experiments, conducted across twelve datasets involving over 4,000,000 molecular pairs, exhibit the superiority of our method over current GNN and LLM-based baselines. Code is available at https://github.com/MangoKiller/MolTC.
@@ 1 Introduction
(2) Dataset Foundation Construction: In sight of the absence of a comprehensive MRL datasets for biochemical LLMs, we construct a Molecular inTeractive instructions dataset, termed MoT-instruction. Specifically, we first conduct twelve well-established MRL datasets across various domain, and source their detailed molecular properties from authoritative biochemical databases. Based on this, we meticulously compile these properties and empirically determine their optimal instructions. These process ensures that MoT-instructions can not only enhance the performance of our MolTC, but also contribute to the development of other biochemical LLMs involving MRL.
@@ 3 Experiment
Baseline Model FreeSolv Abraham CompSol CombiSolv
@@
GNN Based CIGIN $0.589_{\pm 0.053}$ $0.931_{\pm 0.066}$ $0.314_{\pm 0.004}$ $0.607_{\pm 0.011}$ $0.197_{\pm 0.003}$ $0.349_{\pm 0.005}$ $0.288_{\pm 0.005}$ $0.664_{\pm 0.012}$
D-MPNN $0.702_{\pm 0.014}$ $1.231_{\pm 0.029}$ $0.484_{\pm 0.012}$ $0.705_{\pm 0.025}$ $0.205_{\pm 0.006}$ $0.373_{\pm 0.007}$ $0.482_{\pm 0.013}$ $0.895_{\pm 0.055}$
GEM $0.598_{\pm 0.018}$ $1.188_{\pm 0.049}$ $\underline{0.254}_{\pm 0.004}$ $0.531_{\pm 0.005}$ $0.203_{\pm 0.006}$ $0.337_{\pm 0.007}$ $0.290_{\pm 0.009}$ $0.783_{\pm 0.020}$
CGIB $\underline{0.541}_{\pm 0.009}$ $\underline{0.917}_{\pm 0.055}$ $0.258_{\pm 0.008}$ $\underline{0.530}_{\pm 0.009}$ $\underline{0.178}_{\pm 0.004}$ $\underline{0.301}_{\pm 0.003}$ $\underline{0.230}_{\pm 0.004}$ $\underline{0.394}_{\pm 0.009}$
ML Based GOVER $0.636_{\pm 0.026}$ $1.074_{\pm 0.049}$ $0.347_{\pm 0.005}$ $0.625_{\pm 0.016}$ $0.184_{\pm 0.005}$ $0.371_{\pm 0.014}$ $0.412_{\pm 0.016}$ $0.728_{\pm 0.034}$
SolvBert $0.602_{\pm 0.029}$ $1.034_{\pm 0.044}$ ${0.496}_{\pm 0.007}$ ${0.693}_{\pm 0.014}$ $0.192_{\pm 0.008}$ $0.353_{\pm 0.008}$ ${0.418}_{\pm 0.018}$ ${0.711}_{\pm 0.020}$
Uni-Mol ${0.575}_{\pm 0.060}$ ${1.012}_{\pm 0.070}$ $0.355_{\pm 0.007}$ $0.602_{\pm 0.024}$ $0.198_{\pm 0.002}$ $0.344_{\pm 0.003}$ $0.267_{\pm 0.005}$ $0.669_{\pm 0.017}$
SMD $0.599_{\pm 0.037}$ $1.202_{\pm 0.036}$ $0.400_{\pm 0.022}$ $0.646_{\pm 0.037}$ $0.199_{\pm 0.006}$ $0.348_{\pm 0.007}$ $0.657_{\pm 0.011}$ $1.023_{\pm 0.029}$
LLM Based Galactica $0.882_{\pm 0.010}$ $1.438_{\pm 0.066}$ $0.645_{\pm 0.008}$ $1.064_{\pm 0.016}$ $0.594_{\pm 0.006}$ $0.854_{\pm 0.008}$ $0.831_{\pm 0.018}$ $1.486_{\pm 0.035}$
Chem T5 $0.802_{\pm 0.036}$ $1.377_{\pm 0.057}$ $0.629_{\pm 0.010}$ $0.910_{\pm 0.017}$ $0.445_{\pm 0.008}$ $0.734_{\pm 0.010}$ $0.882_{\pm 0.015}$ $1.297_{\pm 0.024}$
MolCA $0.760_{\pm 0.033}$ $1.271_{\pm 0.039}$ $0.581_{\pm 0.007}$ $0.897_{\pm 0.008}$ $0.467_{\pm 0.006}$ $0.716_{\pm 0.022}$ $0.648_{\pm 0.033}$ $1.125_{\pm 0.035}$
MolT5 $0.705_{\pm 0.047}$ $1.135_{\pm 0.069}$ $0.549_{\pm 0.008}$ $0.832_{\pm 0.006}$ $0.476_{\pm 0.003}$ $0.695_{\pm 0.013}$ $0.652_{\pm 0.023}$ $1.124_{\pm 0.027}$
MolTC (Ours) $0.502_{\pm 0.011}$ $0.684_{\pm 0.042}$ $0.194_{\pm 0.009}$ $0.388_{\pm 0.010}$ $0.171_{\pm 0.006}$ $0.295_{\pm 0.004}$ $0.172_{\pm 0.004}$ $0.465_{\pm 0.008}$
@@ 3.1 Experimental Setting
We evaluate MolTC on twelve well-established downstream molecule interaction tasks involving qualitative and quantitative analysis. Here we provide a brief overview of our experimental setup. Detailed descriptions are presented in the appendix.
Datasets. We employ 12 datasets across various domains such as DDI, SSI, and CSI. Specifically, we collect Drugbank (Version 5.0.3), ZhangDDI Zhang et al. (2017), ChChMiner Zitnik et al. (2018), DeepDDI Ryu et al. (2018), TWOSIDES Tatonetti et al. (2012), Chromophore Joung et al. (2020), MNSol Marenich et al. (2020), CompSol Moine et al. (2017), Abraham Grubbs et al. (2010), CombiSolv Vermeire and Green (2021), FreeSolv Mobley and Guthrie (2014) and CombiSolv-QM Vermeire and Green (2021).
@@ 3.3 Quantitative Prediction Performance (RQ2)
Obs.3: MolTC continues to lead in quantitative analysis tasks, an area typically challenging for LLMs. Despite the strong baseline set by CGIB, characterized by low MAE and RMSE across datasets, MolTC outperforms it in every metric. For instance, it achieves a 23.98% reduction in RMSE on the CombiSolv dataset relative to CGIB. This underscores the advantage of adeptly leveraging the interaction between SMILE representations and molecular graph structures.
@@ Appendix A Related Work
Furthermore, it is noteworthy that due to the significance of molecular tasks such as DDI for biochemical LLMs Fang et al. (2023d), the Mol-instructions dataset has been introduced more recently Fang et al. (2023c). This dataset shares many similarities with our work, and can be considered as complementary to each other.
@@ B.1 Datasets
We employ 12 datasets across various domains such as DDI, SSI and CSI.
@@
FreeSolv. Mobley and Guthrie (2014) provides 643 experimental and calculated hydration free energy of small molecules in water. In this work, we consider 560 experimental results following previous work.
@@
SSI MNSol MAE $0.584\pm{0.095}$ $0.504\pm{0.038}$ $0.491\pm{0.053}$ $\underline{0.449}\pm{0.081}$ $0.324\pm{0.019}$
RMSE $1.002\pm{0.101}$ $0.973\pm{0.079}$ $0.930\pm{0.062}$ $\underline{0.858}\pm{0.069}$ $0.585\pm{0.023}$
@@
Table 5: Comparative performance of various methods in CombiSolv-QM. The best-performing methods are highlighted with a gray background, while the second-best methods are underscored for emphasis.
Baseline Model CombiSolv-QM
MAE RMSE
GNN Based CIGIN $0.077_{\pm 0.002}$ $0.176_{\pm 0.004}$
D-MPNN $0.116_{\pm 0.006}$ $0.208_{\pm 0.005}$
GEM $0.079_{\pm 0.003}$ $0.162_{\pm 0.002}$
CGIB $\underline{0.074}_{\pm 0.004}$ $\underline{0.150}_{\pm 0.005}$
ML Based GOVER $0.094_{\pm 0.003}$ $0.277_{\pm 0.005}$
SolvBert $0.102_{\pm 0.005}$ $0.318_{\pm 0.006}$
Uni-Mol ${0.089}_{\pm 0.006}$ ${0.214}_{\pm 0.005}$
SMD $0.107_{\pm 0.004}$ $0.341_{\pm 0.003}$
LLM Based Galactica $0.303_{\pm 0.004}$ $0.601_{\pm 0.008}$
Chem T5 $0.321_{\pm 0.006}$ $0.555_{\pm 0.008}$
MolCA $0.298_{\pm 0.004}$ $0.545_{\pm 0.007}$
MolT5 $0.214_{\pm 0.004}$ $0.339_{\pm 0.009}$
MolTC (Ours) $0.072_{\pm 0.002}$ $0.140_{\pm 0.003}$
@@ B.6 More Experimental Results
Table 4 presents the experimental results not shown in the main text due to length constraints. Note that the three datasets in the CSI domain are all derived by splitting the Chromophore dataset. As discussed in Section 3.3, for a fair comparison, we limited the input features to four-dimensional molecular information, comprising atom type, chirality tag, bond type, and bond direction. Given the difficulty of convergence for some DL-based baselines under this setting, we only showcased the performance of the LLM-based baselines. Meanwhile, considering that our SSI tasks are firstly fine-tuned on the CombiSolv-QM dataset, we present the comprehensive results of this dataset, as shown in Table 5. Observations from Table 4 and 5 are largely consistent with those in the main experimental section. That is, across all tasks, our MolTC outperforms the LLM-based baseline methods in a large margin.
@@ Appendix C Future Work
In this paper, we introduce a novel unified framework, leveraging LLM technology to predict molecular interactive properties. The future development directions of this project are twofold. First, there is an emphasis on expanding its application scope, for instance, applying it to downstream tasks such as few-shot learning. Second, we aim to enhance its capabilities by incorporating technologies like graph explainability Fang et al. (2023a, b), graph sampling Wang et al. (2022, 2023); Fang et al. (2024), and spatio-temporal modeling Xia et al. (2023); Wu et al. (2023), making it more comprehensive or enabling it to process multiple inputs simultaneously, instead of just two.